# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 2.2 Wolff axioms and their generalizations
Let $N$ be the largest integer with $C^{N+1}\leq\delta^{-1}$. For $i=1,\ldots,N$, let $\mathbb{T}_{\tau_{i}},$ be a set of $\tau_{i}$-tubes, with $\tau_{i}\in[\delta C^{i},\delta C^{i+1})$ that satisfy Items (i) and (ii) from Definition 2.12. We claim that for $i<j$, if $T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}},T_{\tau_{j}}\in\mathbb{T}_{\tau_{j}}$, and $\mathbb{T}[T_{\tau_{i}}]\cap\mathbb{T}[T_{\tau_{j}}]\neq\emptyset$, then $\mathbb{T}[T_{\tau_{i}}]\subset\mathbb{T}[T_{\tau_{j}}]$. To see this, observe that if the intersection contains at least one tube $T\in\mathbb{T}$, then $2T_{\tau_{j}}\supset N_{\tau_{j}}(T)\supset T_{\tau_{i}}$. Since $\mathbb{T}_{\tau_{j}}$ covers $\mathbb{T},$ and the sets $\{\mathbb{T}[2T_{\tau_{j}}^{\prime}]\colon T_{\tau_{j}}^{\prime}\in\mathbb{T}_{\tau_{j}}\}$ are disjoint, we conclude that $\mathbb{T}[T_{\tau_{i}}]\subset\mathbb{T}[T_{\tau_{j}}]$.
The consequence of the above observation is the following: if we partially order the sets $\{\mathbb{T}[T_{\tau_{i}}]\colon T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}},\ i=1,\ldots,N\}$ under inclusion, then we actually obtain a total ordering, and this ordering forms a tree with $N$ levels—the vertices at level $i$ are precisely the sets $\mathbb{T}[T_{\tau_{i}}]$. By Item (i) of Definition 2.12, for each $i=1,\ldots,N$ the sets $\{\mathbb{T}[T_{\tau_{i}}]\colon T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}}\}$ have comparable cardinality, up to a multiplicative factor of $C$. This tree is precisely what is needed to verify Lemma 2.13: given $\rho_{0}\in[\delta,1]$, we select an index $i$ so that $\tau_{i}\in[\rho_{0},C\rho_{0})$, and we choose $\rho=\tau_{i}$, $\mathbb{T}_{\rho}=\mathbb{T}_{\tau_{i}}$; note that since $\mathbb{T}_{\rho}$ forms a partitioning cover of $\mathbb{T}$, the $\rho$-tubes in $\mathbb{T}_{\rho}$ are essentially distinct. ∎
@@
Let $C\geq 1$ and let $\mathbb{T}$ be a set of $\delta$-tubes in $\mathbb{R}^{n}$. We say that $\mathbb{T}$ satisfies a Frostman condition at dimension $\sigma$ with error $C$ if the tubes in $\mathbb{T}$ are essentially distinct, and for every $\rho\geq\delta$ and every $\rho$-tube $T_{\rho}$, we have #$\#\mathbb{T}[T_{\rho}]\leq C\rho^{\sigma}(\#\mathbb{T})$.
@@
Let $C\geq 1$ and let $\mathbb{T}$ be a set of $\delta$-tubes in $\mathbb{R}^{n}$. We say that $\mathbb{T}$ satisfies the Self-Similar Convex Wolff Axioms with error $C$ if the tubes in $\mathbb{T}$ are essentially distinct, and for every $\rho_{0}\in[\delta,1]$, there exists $\rho\in[\rho_{0},C\rho_{0})$ and a set of $\rho$-tubes $\mathbb{T}_{\rho}$ that satisfies the following properties.
@@
    For each $T_{\rho}\in\mathbb{T}_{\rho}$, the unit re-scaling of $\mathbb{T}[T_{\rho}]$ relative to $T_{\rho}$ satisfies the Convex Wolff Axioms with error $C$.
@@ 2.3 Discretized Assouad dimension
If the scale $\delta$ is apparent from context, then for berevity we may say “$E$ has discretized Assouad dimension at least $\beta$ at scale separation $A$.”
In this language, the conclusion of Theorem 1.5 says that the set $\bigcup_{T\in\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale separation $\delta^{-\eta}.$
@@ 3 Assouad-extremal Kakeya sets
In this section, we will suppose that Theorem 1.5 is false, and we seek to construct a “worst possible” counter-example to Theorem 1.5. In this and later sections, we will restrict attention to $\mathbb{R}^{3}$. Our example will be worst possible in two respects. First, Inequality (1.1) will fail as dramatically as possible, i.e. a reverse inequality will hold where the term $(\rho/r)^{\varepsilon}$ on the RHS of (1.1) will be replaced by $(\rho/r)^{\omega}$ for the largest possible value of $\omega>0$. Second, of all collections of tubes for which such a reverse inequality holds for this value of $\omega$, we will choose a set $\mathbb{T}$ of largest possible cardinality. The hypotheses of Theorem 1.5 ensure that $\mathbb{T}$ has size at least $\delta^{-2}$; we will choose a set of tubes of size $\delta^{-\alpha}$, for the largest possible value of $\alpha$.
@@ 3.1 Defining $\omega$
Let $(\mathbb{T},Y)_{\delta}$ be a set of tubes and their associated shading. Let $\delta\leq\rho<r\leq 1$, and let $B$ be a ball of radius $r$ whose intersection with $\bigcup Y(T)$ has positive measure. Define $\zeta=\zeta(\mathbb{T},Y,B,\rho)$ to be the unique number satisfying
@@
where the infimum is taken over all pairs $(B,\rho)$, where $\rho\in[\delta,\delta^{\eta}]$ and $B$ is a ball of radius $r\in[\delta^{-\eta}\rho,1]$ whose intersection with $\bigcup Y(T)$ has positive measure. Observe that as $\eta\searrow 0$, the set of admissible pairs $(B,\rho)$ becomes larger, and hence $\zeta(\mathbb{T},Y;\eta)$ weakly decreases, i.e. $\zeta(\mathbb{T},Y;\eta)$ is a weakly increasing function of $\eta$.
@@
Again, $\omega(\varepsilon,\eta,\delta;u)$ is a weakly increasing function of $\eta$. It is also a weakly increasing function of $\varepsilon$ and weakly decreasing function of $u$, since as $\varepsilon$ become larger or $u$ becomes smaller, the set of admissible pairs $(\mathbb{T},Y)_{\delta}$ becomes larger.
@@
Let $\omega\geq 0$ be as in Definition 3.1. For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct tubes that satisfy the Convex Wolff Axioms with error $\delta^{-\eta}$, with a $\delta^{\eta}$-dense shading.
Then $\bigcup_{T\in\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\omega-\varepsilon$ at scale separation $\delta^{-\eta}$.
@@ 3.2 Defining $\alpha$
The quantity $\alpha$ has the following interpretation. We have defined $3-\omega$ to be the smallest possible discretized Assouad dimension of a collection of essentially distinct tubes that satisfies the Convex Wolff Axioms. Every such collection of tubes must have cardinality at least $\delta^{-2}$, and in general we might expect that larger collections of tubes should have larger discretized Assouad dimension. Any collection of essentially distinct tubes of cardinality substantially larger than $\delta^{-\alpha}$ that satisfies the Convex Wolff Axioms must have discretized Assouad dimension larger than $3-\omega$.
@@ 3.3 The existence of Assouad-Extremal Kakeya sets
For $\varepsilon>0,$ we say a pair $(\mathbb{T},Y)_{\delta}$ is $\varepsilon$ Assouad-extremal if it satisfies the following properties.
@@
In light of (3.3), there exists a set of tubes and their associated shading $(\mathbb{T},Y)_{\delta}$, so that (i) #$\#\mathbb{T}\geq\delta^{-u+\varepsilon/2}\geq\delta^{-\alpha+\varepsilon}$; (ii) the tubes are contained in $B(0,1)$, essentially distinct, and satisfy the Convex Wolff Axioms with error $\delta^{-\varepsilon}$; (iii) $Y$ is a $\delta^{\varepsilon}$-dense shading; and (iv) we have $\zeta(\mathbb{T},Y;\eta)\geq\omega-\varepsilon$. Thus by (3.2), for every pair of scales $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\varepsilon}r$ and every ball $B$ of radius $r$, we have
@@
But this is precisely (3.5). ∎
@@ 4 Convex Wolff Axioms and multi-scale structure
The purpose of this section is to prove the following result.
@@
For all $\varepsilon>0$, there exists $\tau>0$, so that for all $\eta_{1}>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\beta\geq 2$ and let $(\mathbb{T},Y)_{\delta}$ be a set of $\delta^{-\beta}$ essentially distinct $\delta$-tubes that satisfy the Convex Wolff Axioms with error $\delta^{\eta}$, with $Y$ a $\delta^{\eta}$ uniformly dense shading. Then at least one of the following must be true.
@@
    $\bigcup Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$, with scale separation $\delta^{-\eta}$.
@@
    There exists $\rho\in[\delta,\delta^{\varepsilon}]$ and a set $\tilde{\mathbb{T}}$ of $\rho$-tubes, with the following properties.
@@
    There exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}[T_{\rho}]$, so that the following holds.
@@
Remark 4.2.
Conclusions (A) and (B) are somewhat self-explanatory. Conclusion (C) says that at a coarser scale, we can find a set of tubes that is richer than our original set $\mathbb{T}$. Conclusion (D) says that we can find a $\rho$-tube so that the unit-rescaling of $\mathbb{T}[T_{\rho}]$ is richer than our original set $\mathbb{T}$.
An important step in proving Proposition 4.1 is to study arrangements of tubes that cluster into rectangular prisms. In the next section, we will show that unions of rectangular prisms of dimensions $s\times t\times 1$ with $s<\!\!<t$ have large Assouad dimension.
@@ 4.1 Arrangements of rectangular prisms
The next result is a straightforward consequence of Cauchy-Schwartz; it says that a union of $\delta\times 1\times 1$ prisms satisfying the Convex Wolff Axioms must have almost maximal volume.
@@
Let $0<\delta\leq 1$ and $K\geq 1$. Let $\mathcal{R}$ be a multi-set of $\delta\times 1\times 1$ rectangular prisms, and suppose that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $K$. Let $\{Y(R),R\in\mathcal{R}\}$ be a $K^{-1}$-dense shading. Then
$\Big{|}\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\gtrsim K^{-3}|\log\delta|^{-1}.$ (4.1)
Proof.
For each prism $R\in\mathcal{R}$, there is an associated normal vector $v(R)\in S^{2}$, which is well-defined up to uncertainty $O(\delta)$. For $R,R^{\prime}\in\mathcal{R}$, define $\angle(R,R^{\prime})$ to be the (unsigned) angle between $v(R)$ and $v(R^{\prime})$.
Fix $R\in\mathcal{R}$. We have
$\sum_{R^{\prime}\in\mathcal{R}}|R\cap R^{\prime}|$ $=\sum_{j=0}^{\log(1/\delta)}\sum_{\begin{subarray}{c}R^{\prime}\colon R\cap R^{\prime}\neq\emptyset,\\ \angle(R,R^{\prime})\sim 2^{j}\delta\end{subarray}}|R\cap R^{\prime}|\lesssim\sum_{j=0}^{\log(1/\delta)}\Big{(}K(2^{j}\delta\#\mathcal{R})\Big{)}\Big{(}\frac{\delta^{2}}{2^{j}\delta}\Big{)}$
$\lesssim K\delta^{2}(\#\mathcal{R})\log(1/\delta).$
In the above computation, observe that if $R\cap R^{\prime}$ is non-empty and $\angle(R,R^{\prime})=\theta\sim 2^{j}\delta$, then $R^{\prime}$ is contained in the $O(\theta)\times 1\times 1$ rectangular prism concentric with $R$. Since $\mathcal{R}$ obeys the Convex Wolff Axioms, at most $K(2^{j}\delta)(\#\mathcal{R})$ slabs $R^{\prime}\in\mathcal{R}$ can have this property. We conclude that
$\Big{\|}\sum_{R\in\mathcal{R}}\chi_{R}\Big{\|}_{2}^{2}\lesssim K\delta^{2}(\#\mathcal{R})^{2}\log(1/\delta).$ (4.2)
$K^{-1}\delta(\#\mathcal{R})\leq\int\chi_{E}\cdot\sum_{R\in\mathcal{R}}\chi_{Y(R)}\leq\int\chi_{E}\cdot\sum_{R\in\mathcal{R}}\chi_{R}\leq|E|^{1/2}\Big{\|}\sum_{R\in\mathcal{R}}\chi_{R}\Big{\|}_{2}.$ (4.3)
Rearranging and using (4.2), we obtain (4.1). ∎
Next, we will establish a geometric argument showing that a union of $s\times t\times 1$ prisms, $s\ll t$, satisfying the Convex Wolff Axioms must either have nearly maximal discretized Assouad dimension, or the prisms must arrange themselves into $s^{\prime}\times t^{\prime}\times 1$ prisms for $s^{\prime}$ and $t^{\prime}$ substantially larger than $s$ and $t$, respectively. The precise statement is as follows.
Lemma 4.4.
For all $\varepsilon,\tau\in(0,1]$, there exists $\delta_{0},\eta>0$ so that the following is true for all $\delta\in(0,\delta_{0}]$. Let $\delta\leq s\leq t\leq 1$ with $s\leq\delta^{\varepsilon}t$. Let $\mathcal{R}$ be a multi-set of $s\times t\times 1$ prisms contained in $B(0,1)$, and suppose that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$. Let $\{Y(R),R\in\mathcal{R}\}$ be a $\delta^{\eta}$-dense shading.
Then at least one of the following two things must happen.
(A): There exists $r\in[\delta^{-\eta}s,1]$ and a $r$-ball $B$ so that
@@
(B): There are scales $s^{\prime}\in[\delta^{-\varepsilon^{2}/2}s,\delta^{\varepsilon/2}]$ and $t^{\prime}=\min\big{(}1,t\frac{s^{\prime}}{s}\big{)}$, and a multi-set $\mathcal{R}^{\prime}$ of $s^{\prime}\times t^{\prime}\times 1$ prisms that satisfies the Convex Wolff Axioms with error $\delta^{-\tau}$, so that the shading
@@
After dyadic pigeonholing, we can select a multiplicity $\mu$ and a $\geq\frac{1}{10}|\log\delta|^{-1}\delta^{\eta}$-dense shading $Y_{1}(R)\subset Y(R)$ so that $\sum_{R\in\mathcal{R}}\chi_{Y_{1}(R)}(x)\sim\mu$ for all $x\in\bigcup_{R\in\mathcal{R}}Y_{1}(R)$. Let $Y_{2}(R)\subset Y_{1}(R)$ be a regular (in the sense of Definition 2.2) refinement of $Y_{1}$, with $|Y_{2}(R)|\geq\frac{1}{2}|Y_{1}(R)|$ for each $R\in\mathcal{R}$.
@@
If $\mu\leq\delta^{-\varepsilon^{2}+3\eta}(st\#\mathcal{R})$, then
$\Big{|}\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\geq\mu^{-1}\sum_{R\in\mathcal{R}}|Y_{1}(R)|\gtrapprox\delta^{\varepsilon^{2}-3\eta}\delta^{\eta}\frac{\sum_{R\in\mathcal{R}}|R|}{st(\#\mathcal{R})}\geq\delta^{\varepsilon^{2}-\eta}\geq\delta^{-\eta}s^{\varepsilon},$

### v2

@@ 2.2 Wolff axioms and their generalizations
Let $N$ be the largest integer with $C^{N+1}\leq\delta^{-1}$. For $i=1,\ldots,N$, let $\mathbb{T}_{\tau_{i}},$ be a set of $\tau_{i}$-tubes, with $\tau_{i}\in[\delta C^{i},\delta C^{i+1})$ that satisfies Items (i) and (ii) from Definition 2.12. We claim that for $i<j$, if $T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}},T_{\tau_{j}}\in\mathbb{T}_{\tau_{j}}$, and $\mathbb{T}[T_{\tau_{i}}]\cap\mathbb{T}[T_{\tau_{j}}]\neq\emptyset$, then $\mathbb{T}[T_{\tau_{i}}]\subset\mathbb{T}[T_{\tau_{j}}]$. To see this, observe that if the intersection contains at least one tube $T\in\mathbb{T}$, then $2T_{\tau_{j}}\supset N_{\tau_{j}}(T)\supset T_{\tau_{i}}$. Since $\mathbb{T}_{\tau_{j}}$ covers $\mathbb{T},$ and the sets $\{\mathbb{T}[2T_{\tau_{j}}^{\prime}]\colon T_{\tau_{j}}^{\prime}\in\mathbb{T}_{\tau_{j}}\}$ are disjoint, we conclude that $\mathbb{T}[T_{\tau_{i}}]\subset\mathbb{T}[T_{\tau_{j}}]$.
The consequence of the above observation is the following: if we partially order the sets $\{\mathbb{T}[T_{\tau_{i}}]\colon T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}},\ i=1,\ldots,N\}$ under inclusion, then this forms a tree with $N$ levels—the vertices at level $i$ are precisely the sets $\mathbb{T}[T_{\tau_{i}}]$. By Item (i) of Definition 2.12, for each $i=1,\ldots,N$ the sets $\{\mathbb{T}[T_{\tau_{i}}]\colon T_{\tau_{i}}\in\mathbb{T}_{\tau_{i}}\}$ have comparable cardinality, up to a multiplicative factor of $C$. This tree is precisely what is needed to verify Lemma 2.13: given $\rho_{0}\in[\delta,1]$, we select an index $i$ so that $\tau_{i}\in[\rho_{0},C\rho_{0})$, and we choose $\rho=\tau_{i}$, $\mathbb{T}_{\rho}=\mathbb{T}_{\tau_{i}}$; note that since $\mathbb{T}_{\rho}$ forms a partitioning cover of $\mathbb{T}$, the $\rho$-tubes in $\mathbb{T}_{\rho}$ are essentially distinct. ∎
@@
[…]
@@
Let $C\geq 1$ and let $\mathbb{T}$ be a set of $\delta$-tubes in $\mathbb{R}^{n}$. We say that $\mathbb{T}$ satisfies the self-similar Convex Wolff Axioms with error $C$ if the tubes in $\mathbb{T}$ are essentially distinct, and for every $\rho_{0}\in[\delta,1]$, there exists $\rho\in[\rho_{0},C\rho_{0})$ and a set of $\rho$-tubes $\mathbb{T}_{\rho}$ that satisfies the following properties:
@@
    For each $T_{\rho}\in\mathbb{T}_{\rho}$, the unit rescaling of $\mathbb{T}[T_{\rho}]$ relative to $T_{\rho}$ satisfies the Convex Wolff Axioms with error $C$.
@@ 2.3 Discretized Assouad dimension
If the scale $\delta$ is apparent from context, then for brevity we may say “$E$ has discretized Assouad dimension at least $\beta$ at scale separation $A$.”
In this language, the conclusion of Theorem 1.5 says that the set $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale separation $\delta^{-\eta}.$
@@ 3 Assouad-extremal Kakeya sets
In this section we will suppose that Theorem 1.5 is false, and we seek to construct a “worst possible” counter-example to Theorem 1.5. In this and later sections, we will restrict attention to $\mathbb{R}^{3}$. Our example will be worst possible in two respects. First, Inequality (1.1) will fail as dramatically as possible, i.e. a reverse inequality will hold where the term $(\rho/r)^{\varepsilon}$ on the RHS of (1.1) will be replaced by $(\rho/r)^{\omega}$ for the largest possible value of $\omega>0$. Second, of all collections of tubes for which such a reverse inequality holds for this value of $\omega$, we will choose a set $\mathbb{T}$ of largest possible cardinality. The hypotheses of Theorem 1.5 ensure that $\mathbb{T}$ has size at least $\delta^{-2}$; we will choose a set of tubes of size $\delta^{-\alpha}$, for the largest possible value of $\alpha$.
@@ 3.1 Defining $\omega$
Let $(\mathbb{T},Y)_{\delta}$ be a set of tubes and their associated shading. Let $\delta\leq\rho<r\leq 1$, and let $B$ be a ball of radius $r$ whose intersection with $\bigcup_{\mathbb{T}}Y(T)$ has positive measure. Define $\zeta=\zeta(\mathbb{T},Y,B,\rho)$ to be the unique number satisfying
@@
where the infimum is taken over all pairs $(B,\rho)$, where $\rho\in[\delta,\delta^{\eta}]$ and $B$ is a ball of radius $r\in[\delta^{-\eta}\rho,1]$ whose intersection with $\bigcup_{\mathbb{T}}Y(T)$ has positive measure. Observe that as $\eta\searrow 0$, the set of admissible pairs $(B,\rho)$ becomes larger, and hence $\zeta(\mathbb{T},Y;\eta)$ weakly decreases, i.e. $\zeta(\mathbb{T},Y;\eta)$ is a weakly increasing function of $\eta$.
@@
Again, $\omega(\varepsilon,\eta,\delta;u)$ is a weakly increasing function of $\eta$. It is also a weakly increasing function of $\varepsilon$ and a weakly decreasing function of $u$, since as $\varepsilon$ becomes larger or $u$ becomes smaller, the set of admissible pairs $(\mathbb{T},Y)_{\delta}$ becomes larger.
@@
Let $\omega\geq 0$ be as in Definition 3.1. For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be a set of essentially distinct tubes that satisfy the Convex Wolff Axioms with error $\delta^{-\eta}$, with $Y$ a $\delta^{\eta}$-dense shading.
Then $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\omega-\varepsilon$ at scale separation $\delta^{-\eta}$.
@@ 3.2 Defining $\alpha$
The quantity $\alpha$ has the following interpretation. We have defined $3-\omega$ to be the smallest possible discretized Assouad dimension of a collection of essentially distinct tubes that satisfy the Convex Wolff Axioms. Every such collection of tubes must have cardinality at least $\delta^{-2}$, and in general we might expect that larger collections of tubes should have larger discretized Assouad dimension. Any collection of essentially distinct tubes of cardinality substantially larger than $\delta^{-\alpha}$ that satisfies the Convex Wolff Axioms must have discretized Assouad dimension larger than $3-\omega$.
@@ 3.3 The existence of Assouad-extremal Kakeya sets
For $\varepsilon>0,$ we say a pair $(\mathbb{T},Y)_{\delta}$ is $\varepsilon$ Assouad-extremal if it satisfies the following properties:
  • (i)
@@
In light of (3.3), there exists a set of tubes and their associated shading $(\mathbb{T},Y)_{\delta}$, so that the following holds:
  • (a)
    #$\#\mathbb{T}\geq\delta^{-u+\varepsilon/2}\geq\delta^{-\alpha+\varepsilon}$ (cf. Definition 3.3, Item (i)).
  • (b)
    The tubes are contained in $B(0,1)$, essentially distinct, and satisfy the Convex Wolff Axioms with error $\delta^{-\varepsilon}$ (cf. Definition 3.3, Item (ii)).
  • (c)
    $Y$ is a $\delta^{\varepsilon}$-dense shading (cf. Definition 3.3, Item (iii)).
  • (d)
    We have $\zeta(\mathbb{T},Y;\eta)\geq\omega-\varepsilon$.
Thus by (3.2), for every pair of scales $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\varepsilon}r$ and every ball $B$ of radius $r$, we have
@@
But this is precisely (3.5), and thus $(\mathbb{T},Y)_{\delta}$ satisfies Definition 3.3, Item (iv). ∎
@@ 4 Convex Wolff Axioms and multi-scale structure
In this section we will carefully execute the arguments described in Section 1.1. The main result is Proposition 4.1, which says that under suitable hypotheses, for every collection of $\delta$-tubes $\mathbb{T}$, at least one of the following must occur:
  • (A)
    $\mathbb{T}$ has discretized Assouad dimension close to 3 (as we will see later in this section, this is forced to happen if the tubes in $\mathbb{T}$ arrange themselves into planks).
  • (B)
    $\mathbb{T}$ satisfies an analogue of stickiness.
  • (C)
    A suitable thickening of $\mathbb{T}$ yields a richer collection of tubes.
  • (D)
    A suitable zooming-in and rescaling of $\mathbb{T}$ yields a richer collection of tubes.
The precise statement is as follows.
@@
For all $\varepsilon>0$, there exists $\tau>0$, so that for all $\eta_{1}>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\beta\geq 2$ and let $(\mathbb{T},Y)_{\delta}$ be a set of $\delta^{-\beta}$ essentially distinct $\delta$-tubes that satisfy the Convex Wolff Axioms with error $\delta^{\eta}$, with $Y$ a $\delta^{\eta}$-uniformly dense shading. Then at least one of the following must hold:
@@
    $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$, with scale separation $\delta^{-\eta}$.
@@
    There exists $\rho\in[\delta,\delta^{\varepsilon}]$ and a set $\tilde{\mathbb{T}}$ of $\rho$-tubes, with the following properties:
@@
    There exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}[T_{\rho}]$, so that the following holds:
@@
To prove Proposition 4.1, we will consider arrangements $\mathbb{T}$ for which conclusions (B) and (C) fail. This will force the existence of a $\rho$-tube $T_{\rho}$ that satisfies Item (D.i). However, it is not clear that this $\rho$-tube will also satisfy Item (D.ii); the main concern is that there might exist many convex sets $W\subset T_{\rho}$ of dimensions $a\times b\times 1$ with $a<\!\!<b$ that contain an illegally large number of tubes from $\mathbb{T}[T_{\rho}]$ (convex sets of dimensions $a\times b\times 1$ with $a\sim b$ can be handled by replacing $\rho$ with a smaller scale). Our goal is to show that if this happens, then Conclusion (A) holds. To do this, we will show that if this happens, then the tubes in $\mathbb{T}[T_{\rho}]$ must cluster into planks; this step is straightforward and will be deferred to later. The second step is to show that under suitable hypotheses, every arrangement of planks must have discretized Assouad dimension close to 3; we will do this in the next sub-section.
@@ 4.1 Arrangements of rectangular prisms
We begin the process described at the end of the previous section. Let $\mathcal{R}$ be a set of planks, i.e. rectangular prisms of dimensions roughly $s\times t\times 1$ for some $\delta\leq s\leq t\leq 1$. If a typical pair of planks from $\mathcal{R}$ intersects tangentially, then this forces the planks to cluster into wider (and somewhat thicker) planks; crucially, the aspect ratio (i.e. thickness vs. width) remains the same. If instead a typical pair of planks does not intersect tangentially, then the hairbrush of a typical plank $R\in\mathcal{R}$ (i.e. the union of the other planks that intersect $R$) fills out a thickened neighbourhood of $R$, and this in turn means that the union of planks has discretized Assouad dimension close to 3. The precise statement is as follows.
@@
For all $\varepsilon,\tau\in(0,1]$, there exists $\delta_{0},\eta>0$ so that the following is true for all $\delta\in(0,\delta_{0}]$. Let $\delta\leq s\leq t\leq 1$ with $s\leq\delta^{\varepsilon}t$. Let $\mathcal{R}$ be a multi-set of $s\times t\times 1$ rectangular prisms contained in $B(0,1)$, and suppose that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$. Let $\{Y(R),R\in\mathcal{R}\}$ be a $\delta^{\eta}$-dense shading.
Then at least one of the following two things must happen:
  • (A)
    $\bigcup_{\mathcal{R}}Y(R)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $s$, with scale separation $\delta^{-\eta}$. This means that there exists $r\in[\delta^{-\eta}s,1]$ and a $r$-ball $B$ so that
@@
  • (B)
    There are scales $s^{\prime}\in[\delta^{-\varepsilon^{2}/4}s,\delta^{\varepsilon/2}]$ and $t^{\prime}=\min\big{(}1,t\frac{s^{\prime}}{s}\big{)}$, and a multi-set $\mathcal{R}^{\prime}$ of $s^{\prime}\times t^{\prime}\times 1$ prisms that satisfies the Convex Wolff Axioms with error $\delta^{-\tau}$, so that the shading
@@
Step 1: Robust broadness. After replacing $\mathcal{R}$ by a subset of cardinality $\gtrsim\delta^{\eta}(\#\mathcal{R})$ if necessary, we may suppose that $|Y(R)|\geq\frac{1}{2}\delta^{\eta}|R|$ for each $R\in\mathcal{R}$. Let $Y_{1}(R)\subset Y(R)$ be a regular (in the sense of Definition 2.2) refinement of $Y$, with $|Y_{1}(R)|\geq\frac{1}{2}|Y(R)|$ for each $R\in\mathcal{R}$.
Next, we perform a “two-ends” style robust broadness reduction. Let $\varepsilon_{1}>0$ be a small quantity to be chosen below. For each $x\in\bigcup_{R\in\mathcal{R}}Y_{1}(R)$, let $u_{x}\in[t,1]$ and let $T_{u_{x}}$ be a $u_{x}$-tube that comes within a multiplicative factor of 2 of maximizing the quantity
$u_{x}^{-\varepsilon_{1}}\#\{R\in\mathcal{R}\colon x\in Y_{1}(R),\ R\subset T_{u_{x}}\}.$
Since there exists a 1-tube that contains a $\gtrsim 1$ fraction of $\mathcal{R}$, and since $u_{x}\geq\delta$, we have
#$\#\{R\in\mathcal{R}\colon x\in Y_{1}(R),\ R\subset T_{u_{x}}\}\gtrsim\delta^{\varepsilon_{1}}\#\{R\in\mathcal{R}\colon x\in Y_{1}(R)\}.$
Define the shading
$Y_{2}(R)=\{x\in Y_{1}(R)\colon R\subset T_{u_{x}}\}.$
@@
Then $\sum_{R\in\mathcal{R}}|Y_{2}(R)|\gtrsim\delta^{\varepsilon_{1}}\sum_{R\in\mathcal{R}}|Y_{1}(R)|$. For each $x\in\bigcup_{R\in\mathcal{R}}Y_{2}(R)$, each $t\leq r\leq u_{x}$, and each $r$-tube $T_{r}$, we have
#$\#\{R\in\mathcal{R}\colon x\in Y_{2}(R),\ R\subset T_{r}\}\leq 2(r/u_{x})^{\varepsilon_{1}}\#\{R\in\mathcal{R}\colon x\in Y_{2}(R)\}.$ (4.3)
By dyadic pigeonholing, we can select a multiplicity $\mu$ and a number $t\leq u\leq 1$ so that if we define
$Y_{3}(R)=Y_{2}(R)\cap\Big{\{}x\in\mathbb{R}^{3}\colon\sum_{R\in\mathcal{R}}\chi_{Y_{2}(R)}(x)\in[\mu,2\mu),\ u_{x}\in[u/2,u)\Big{\}},$
then provided we choose $\delta_{0}$ sufficiently small and $\eta<\varepsilon_{1}/2$, we have $\sum_{R\in\mathcal{R}}|Y_{3}(R)|\geq\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}}|R|$.
We first consider the case where $\mu\leq\delta^{-\varepsilon^{2}+2\varepsilon_{1}}(st\#\mathcal{R})$. Then
$\Big{|}\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\geq\mu^{-1}\sum_{R\in\mathcal{R}}|Y_{3}(R)|\geq\delta^{\varepsilon^{2}-2\varepsilon_{1}}\delta^{2\varepsilon_{1}}\frac{\sum_{R\in\mathcal{R}}|R|}{st(\#\mathcal{R})}\geq\delta^{\varepsilon^{2}}\geq s^{\varepsilon},$